The completion of a certain chemical reaction is expressed by the equation y=12+4x-2(x*x). Y is the number of seconds needed to complete the reaction and x is the temperature in degrees CCelsius at which the reaction occurs. If the reaction is complete in 45 seconds, what is the temperature at which the reaction occurs?
Answer by ikleyn(52754)

.

The equation


    12 + 4x - 2x^2 = 45


Try to solve it.


    2x^2 - 4x + 33 = 0


    Discriminant  d = b^2 - 4ac = 4^2 - 4*2*33 = -248.


Discriminant is negative, so the equation has no solutions.


It looks like the problem is a   F A K E.